# Euclid Quick Data Release (Q1)

### v1

@@
The Euclid mission aims to survey around 14 0001400014\,00014 000 $\deg^{2}$ of extragalactic sky, providing around $10^{5}$ gravitational lens images. Modelling of gravitational lenses is fundamental to estimate the total mass of the lens galaxy, along with its dark matter content. Traditional modelling of gravitational lenses is computationally intensive and requires manual input. In this paper, we use a Bayesian neural network, LEns MOdelling with Neural networks (LEMON), for modelling Euclid gravitational lenses with a singular isothermal ellipsoid mass profile. Our method estimates key lens mass profile parameters, such as the Einstein radius, while also predicting the light parameters of foreground galaxies and their uncertainties. We validate LEMON’s performance on both mock Euclid data sets, real Euclidised lenses observed with Hubble Space Telescope (hereafter HST), and real Euclid lenses found in the Perseus ERO field, demonstrating the ability of LEMON to predict parameters of both simulated and real lenses. Results show promising accuracy and reliability in predicting the Einstein radius, axis ratio, position angle, effective radius, Sérsic index, and lens magnitude for simulated lens galaxies. The application to real data, including the latest Quick Release 1 strong lens candidates, provides encouraging results, particularly for the Einstein radius. We also verified that LEMON has the potential to accelerate traditional modelling methods, by giving to the classical optimiser the LEMON predictions as starting points, resulting in a speed-up of up to 26 times the original time needed to model a sample of gravitational lenses, a result that would be impossible with randomly initialised guesses. This work represents a significant step towards efficient, automated gravitational lens modelling, which is crucial for handling the large data volumes expected from Euclid.
@@ 1 Introduction
Gravitational strong lensing is a rare astrophysical phenomenon that occurs when a massive object, such as a galaxy, bends the light coming from a background distant object. This effect is a prediction of Einstein’s general relativity, and is an important tool in the study of galaxy evolution, given that a gravitational lens can magnify distant galaxies, allowing for the detailed study of structures in galaxies at high redshift (Coe et al. 2013). Moreover, the distortion of the background object gives information on the properties of the deflector. In particular, when the redshifts of source and lens are known, galaxy-galaxy strong lensing allows us to determine with extreme precision the projected total mass within the Einstein radius of the foreground galaxy, without requiring any modelling a priori. By assuming a model for the deflector mass profile, it is possible to infer more quantities related to the dark matter content of galaxies, such as the total dark matter mass and the dark matter fraction (Koopmans et al. 2006; Tortora et al. 2010; Auger et al. 2010; Sengül & Dvorkin 2022), along with other properties, such as the mass density slope and the initial mass function (Gavazzi et al. 2007; Koopmans et al. 2009; Treu et al. 2010; Sonnenfeld et al. 2013; Shajib et al. 2021); finally, through gravitational lensing, it is also possible to detect substructures of the lens galaxy (Vegetti et al. 2010). Galaxy-galaxy strong lensing is also a powerful probe of the small-scale structure of galaxy clusters (e.g., Meneghetti et al. 2020, 2022, 2023).
The Euclid mission will observe around 14 0001400014\,00014 000 $\deg^{2}$ of extragalactic sky from the Sun-Earth Lagrange point L2, surveying of order $10^{9}$ galaxies (for a description of the Euclid Wide Survey, we refer the reader to Euclid Collaboration: Scaramella et al. 2022; Euclid Collaboration: Mellier et al. 2024, while an overall description of the mission can be found in Laureijs et al. 2011).
@@
Following Euclid’s launch in 2023, ESA and collaborators initiated the Early Release Observations programme (ERO, Euclid Early Release Observations 2024; Cuillandre et al. 2024a), targeting 17 astronomical objects, including galaxy clusters, nearby galaxies, globular clusters, and star-forming regions, with the objective of demonstrating the telescope’s capabilities and providing early scientific insights (Atek et al. 2024; Cuillandre et al. 2024b; Hunt et al. 2024; Kluge et al. 2024; Marleau et al. 2024; Massari et al. 2024; Saifollahi et al. 2024). Among these targets, the Perseus cluster field is especially of interest for strong lensing science, with two main projects underway: the ERO Lens Search Experiment (ELSE), where a blind visual search of galaxy-scale strong lensing systems is carried out (Acevedo Barroso et al. 2024), finding 16 lens candidates (of which five have a convincing lens model) and a parallel search of strong lenses via the use of neural networks (Pearce-Casey et al. 2024) aimed at testing their performance on a real Euclid field. A similar search with neural networks in the other ERO fields is carried out in Nagam et al. (2025). The first statistically relevant sample of around $500$ strong lenses in Euclid has finally been collected with the data from the Euclid Quick Data Release 1 (Q1, Euclid Quick Release Q1 2025; Euclid Collaboration: Aussel et al. 2025), which also represents the first publicly available data set of Euclid Wide Survey (EWS)-like images, processed with the same pipeline and reaching the same depth as the final survey. Despite covering only $63.1$ $\deg^{2}$, Q1 provides a powerful testbed for Euclid’s lens searching capabilities, with the strong lensing discovery engine at the forefront of these efforts (Euclid Collaboration: Walmsley et al. 2025; Euclid Collaboration: Rojas et al. 2025; Euclid Collaboration: Lines et al. 2025; Euclid Collaboration: Li et al. 2025; Euclid Collaboration: Holloway et al. 2025).
@@
High-resolution imaging, together with a large field of view, are crucial for detecting a large number of gravitational lenses. However, the substantial amount of data coming from the mission will necessarily require automated methods for both lens finding and lens modelling. Currently, convolutional neural networks (CNNs) are already used extensively in lens-finding tasks, having generally good performance (Petrillo et al. 2017, 2019a, 2019b; Metcalf et al. 2019; Cañameras et al. 2020; Li et al. 2020; Rezaei et al. 2022; Rojas et al. 2022; Nagam et al. 2024). Instead, modelling of gravitational lenses has traditionally relied on complex and time-consuming techniques that require manual inputs, such as maximum likelihood and Markov chain Monte Carlo methods (Keeton 2016; Nightingale et al. 2021a, b; Harvey-Hawes & Wiltshire 2024). The advent of machine learning has allowed for the automation and acceleration of this process. In particular, CNNs and Bayesian neural networks (BNNs) have shown remarkable results in recovering lens parameters (Perreault Levasseur et al. 2017; Pearson et al. 2019, 2021; Schuldt et al. 2021, 2023; Gentile et al. 2023).
@@
The main algorithm for this work is LEns MOdelling with Neural networks (LEMON; Gentile et al. 2023), a BNN able to perform fast automated analysis of strong gravitational lenses. In the original work, Gentile et al. (2023) used LEMON to estimate three parameters of a strong gravitational lens modelled as a singular isothermal ellipsoid (SIE), the Einstein radius and the two components of the ellipticity, for mock HST and Euclid gravitational lenses. Along with the expected values, LEMON is also able to output uncertainties for each parameter.
@@
In this work, we expand the results of Gentile et al. (2023), by predicting both mass and light parameters of the foreground galaxies in isolated galaxy-galaxy gravitational lenses, considering a single Sérsic model for the light profile and using mock Euclid lenses with contaminants in the image for both training and testing. We will also verify the behaviour of both uncertainty components and verify the capability of LEMON to generalise to real Euclidised HST lens images.
@@
The paper is structured as follows. In Sect. 2 we give details on the various data sets used for the analysis. In Sect. 3 we briefly introduce the LEMON algorithm. In Sect. 4 we describe the training procedure, the metrics used to quantify the performance of LEMON, and the calibration procedure for the total uncertainty. Section 5 is dedicated to the results of the analysis on the simulated test set, while Sect. 6 shows the results associated with real Euclidised lenses and real Euclid lenses found in the Perseus ERO and Q1 fields. Section 7 shows how the joint use of LEMON and classical modelling methods could provide a speed up of the latter. We finally give our conclusions in Sect. 8.
@@ 2 Data
In the following, we will give details on the various data sets used for the analysis. In Sect. 2.1, the Euclid mock lenses used for training and testing are described. We then describe the real Euclidised lenses used for testing LEMON performance on real lenses in Sect. 2.2.
@@ 2.1 Euclid mock lenses
Table 1: List of parameter ranges for the Euclid mock lenses.
@@
Axis ratio $q$ $[0.11,1.00]$
[…]
@@
Lens Sérsic index $n_{\mathrm{lens}}$ $[2.00,8.00]$
Lens magnitude $m_{\mathrm{lens}}$ $[13.88,24.68]$
 [^1] Only position angle and Sérsic index are extracted from a uniform distribution. This choice has been made to avoid introducing a prior in the training of the network. All the other parameters are extracted from non-analytic, empirical distributions derived from the Flagship simulation(Euclid Collaboration: Castander et al. 2024). The coverage of the parameters in the reported intervals is thus non-uniform.
@@
Figure 1: Distribution of mass and light parameters for the Euclid mock lenses. Blue bars show the distribution of lenses without contaminants in the image, while orange bars show the distribution of lenses with contaminants.
@@
For training and performance-testing of LEMON, we have made joint use of 50 0005000050\,00050 000 lenses ‘with companions’ (i.e., having contaminants other than the main lens in the image), each of which contributes to the projected lensing potential, and 50 0005000050\,00050 000 lenses ‘without companions’. Both sets are simulated by the Strong Lensing Science Working Group of the Euclid Consortium (Metcalf et al., in prep.). The lens images are made by simulating the Euclid visible instrument (VIS, Euclid Collaboration: Cropper et al. 2024) camera output, at the depth of the Euclid Wide Survey. The two samples are selected to capture the full range of environmental scenarios that may occur when observing a gravitational lens, including objects along the line of sight as well as companion or background objects surrounding the lens system. Separating the two sets allows us to also test whether the lack of training on images with companions affects the recovery of the parameters, given that the presence of companions in the image could be mistaken by the network as part of the lensing system (see App. A).
@@
The lenses are constructed by first considering galaxies in the Euclid Flagship simulation (Euclid Collaboration: Castander et al. 2024), taken from a sky area of the order of $100\,\textrm{deg}^{2}$, under the requirements that the galaxy chosen is brighter than the lens galaxy magnitude limit of 24 in the $I_{\scriptscriptstyle\rm E}$ band, which is the VIS imaging band, and categorised as a central galaxy. Central galaxies are galaxies that are positioned at the centres of the dark matter halos. Large halos also contain satellite galaxies, which should not produce isolated lenses like the ones we wish to model here. For each selected candidate, the apparent magnitude, effective radius, position angle, axis ratio and redshift are taken from the Flagship and converted into a surface brightness model. In general, the light model is composed of a disc and a spheroid component, but in our catalogue, only galaxies with a spheroid component are considered. This spheroid component is represented by a single Sérsic profile. The Sérsic index is uniformly distributed between 2 and 8. A mass model is then chosen by considering analytical profiles centred on each galaxy. In our catalogue, all lens mass profiles are considered to be SIE profiles.
@@
The mass in the lens is normalised with the ratio of dark matter to stellar mass within one effective radius. The stellar mass and effective radius are taken from the Flagship simulation. The dark matter fraction within one effective radius is taken to be normally distributed with a mean of 0.6 and a standard deviation of 0.1 consistent with observations (Mukherjee et al. 2022). The axis ratio and position angle of the mass is set to those of the light.
@@
Once the lens galaxies have been selected, a source galaxy is placed in the relative background. This increases the sample size of lenses in the data set, in order to offset the inherent rarity of ‘natural lenses’ (i.e. lenses in the Flagship simulation that naturally have a source behind them when the image is traced onto the observer plane). The redshift of the source is sampled randomly from the following distribution:
@@
where $\gamma=-0.23$ and $z_{0}=3.06$ are values found by fitting to the COSMOS 2020 Farmer catalogue $I$-band number counts (Weaver et al. 2022). The surface brightness is represented by between one to four Sérsic profiles. To emulate some of the complex morphologies of high-redshift sources, the number of Sérsic profiles, their relative brightnesses, and their positions are generated randomly from distributions tuned by eye. Their total brightnesses and overall sizes are based on randomly selected sources from the Hubble Ultra Deep Field (HUDF) with similar redshifts (see Meneghetti et al. 2008, 2010 for a discussion of this sample).
Ray-shooting through the light cones is performed, to construct an image of the lens. The deflection caused by other objects along the line of sight is not included, so that the mass distribution remains an SIE in all cases. The lens is then rejected or accepted based on an observability criterion. This criterion seeks to filter out lenses that would not be recognisable as lenses. It combines the signal-to-noise ratio (S/N) in the lensed images, the contrast between the source and lens surface brightnesses and the morphology of the lensed source images. More details on the procedure are available in Metcalf et al. (in prep.).
[…]
The six parameters listed in Table 1 are the main ones that we want to predict with LEMON (the recovery of a seventh parameter, the Einstein mass, is discussed in App. B). They are listed as follows.
@@
    where $A_{\textrm{Ein}}$ is the Einstein area in square radians, defined as the area of the largest critical curve, and the factor 648 000648000648\,000648 000 corresponds to converting $\pi$ radians into arcseconds.
@@
    Axis ratio, defined as $q=b/a$, where $b$ and $a$ are, respectively, the semi-minor and semi-major axes of the ellipsoidal distribution.
@@
    Position angle, defined as the clockwise angle of the semi-major axis of the ellipsoidal distribution, starting from left to top with respect to the image, in units of degrees.
@@
    Circularised effective radius of the lens $R_{\textrm{e}}$, i.e., the radius enclosing one-half of the lens galaxy light, obtained from the major axis effective radius $R_{\textrm{e, maj}}$ via the formula $R_{\textrm{e}}=\sqrt{q}\,R_{\textrm{e, maj}}$. We will refer to it as $R_{\textrm{e, lens}}$ in the following.
@@
    Lens Sérsic index, which regulates the slope of the surface brightness profile. It corresponds to the parameter $n$ that appears in the Sérsic profile definition
@@
    where $I_{\textrm{e}}=I(r=R_{\textrm{e}})$ and $b_{n}$ is a parameter, solution to the equation $\Gamma(2n)=2\,\gamma(2n,b_{n})$, where $\Gamma$ and $\gamma$ are the complete and incomplete gamma functions, respectively. We will refer to it as $n_{\textrm{lens}}$ in the following.
@@
    Lens magnitude, defined as the apparent magnitude in the $I_{\scriptscriptstyle\rm E}$ band of the main lens.
It should be noted that, in these simulations, we implicitly assume that the ellipticity and orientation of the mass profile of the lens follows those of the light profile. This assumption generally holds for early-type galaxies, but outliers can be found in real scenarios, mainly due to the presence of faint discs or external galaxies that induce a substantial external shear (see e.g., Gavazzi et al. 2012). This assumption can thus lead to systematic errors in predicting lens parameters of systems where the mass does not follow the light profile, because the network tends to learn to estimate the parameters from the light distribution, rather than from the lensed image of the source, as noted in Pearson et al. (2019). In future works, we also plan to train on lenses for which this assumption does not hold, as was done in Gentile et al. (2023), albeit with more simplistic mock lenses. The original simulated images have dimensions of $20\arcsec\times 20\arcsec$, as shown in Fig. 2, but during the training of LEMON we create $10\arcsec\times 10\arcsec$ cutouts of these images, centred on the lenses, in order to better see the lens features.
Figure 2: $20\arcsec\times 20\arcsec$ original VIS images of some examples of Euclid mock lenses used for training LEMON. Top row: lenses with companions in the image. Bottom row: lenses without companions.

### v2

@@
The Euclid mission aims to survey around 14 00014\,000 $\deg^{2}$ of extragalactic sky, providing around $10^{5}$ gravitational lens images. Modelling of gravitational lenses is fundamental to estimate the total mass of the lens galaxy, along with its dark matter content. Traditional modelling of gravitational lenses is computationally intensive and requires manual input. In this paper, we use a Bayesian neural network, LEns MOdelling with Neural networks (LEMON), to model Euclid gravitational lenses with a singular isothermal ellipsoid mass profile. Our method estimates key lens mass profile parameters, such as the Einstein radius, while also predicting the light parameters of foreground galaxies and their uncertainties. We validate LEMON’s performance on both mock Euclid datasets, real lenses observed with Hubble Space Telescope (HST) that have been degraded to match observations with the same depth of the Euclid Wide Survey, and real Euclid lenses, demonstrating the ability of LEMON to predict parameters of both simulated and real lenses. Results show promising accuracy and reliability in predicting the Einstein radius, mass and light ellipticities, effective radius, Srsic index, lens magnitude, and unlensed source position for simulated lens galaxies. The application to real data, including the latest Quick Release 1 strong lens candidates, provides encouraging results in the recovery of the parameters for real lenses. We also verified that LEMON has the potential to accelerate traditional modelling methods, by giving to the classical optimiser the LEMON predictions as starting points, resulting in a speed-up of up to 26 times the original time needed to model a sample of gravitational lenses, a result that would be impossible with randomly initialised guesses. Moreover, LEMON can be used to cross-validate results from the traditional modelling methods, and thus has the potential to reduce the failure rate of the Euclid modelling pipeline. This work represents a significant step towards efficient, automated gravitational lens modelling, which is crucial for handling the large data volumes expected from Euclid.
@@ 1 Introduction
Gravitational strong lensing is a rare astrophysical phenomenon that occurs when a massive object, such as a galaxy, bends the light coming from a background distant object. This effect is a prediction of Einstein’s general relativity, and is an important tool in the study of galaxy evolution, given that a gravitational lens can magnify distant galaxies, allowing for the detailed study of structures in galaxies at high redshift (Coe et al. 2013). Moreover, the distortion of the background object gives information on the properties of the deflector. In particular, when the redshifts of source and lens are known, galaxy-galaxy strong lensing allows us to determine with extreme precision the projected total mass within the Einstein radius of the foreground galaxy, without requiring any modelling a priori. By assuming a model for the deflector mass profile, it is possible to infer more quantities related to the dark matter content of galaxies, such as the total dark matter mass and the dark matter fraction (Koopmans et al. 2006; Tortora et al. 2010; Auger et al. 2010; Sengül & Dvorkin 2022), along with other properties, such as the mass density slope and the initial mass function (Gavazzi et al. 2007; Koopmans et al. 2009; Treu et al. 2010; Sonnenfeld et al. 2013; Shajib et al. 2021); finally, through gravitational lensing, it is also possible to detect substructures of the lens galaxy (Vegetti et al. 2010). Galaxy-galaxy strong lensing is also a powerful probe of the small-scale structure of galaxy clusters (e.g. Meneghetti et al. 2020, 2022, 2023).
@@
The Euclid mission will observe around 14 00014\,000 $\deg^{2}$ of extragalactic sky from the Sun-Earth Lagrange point L2, surveying of the order of $10^{9}$ galaxies (for a description of the Euclid Wide Survey, we refer the reader to Euclid Collaboration: Scaramella et al. 2022; Euclid Collaboration: Mellier et al. 2025, while an overall description of the mission can be found in Laureijs et al. 2011). Following Euclid’s launch in 2023, ESA and collaborators initiated the Early Release Observations programme (ERO, Euclid Early Release Observations 2024; Cuillandre et al. 2025a), targeting 17 astronomical objects, including galaxy clusters, nearby galaxies, globular clusters, and star-forming regions, with the objective of demonstrating the telescope’s capabilities and providing early scientific insights (Atek et al. 2025; Cuillandre et al. 2025b; Hunt et al. 2025; Kluge et al. 2025; Marleau et al. 2025; Massari et al. 2025; Saifollahi et al. 2025). Among these targets, the Perseus cluster field is especially of interest for strong lensing science, with two main projects underway: the ERO Lens Search Experiment (ELSE), in which a blind visual search of galaxy-scale strong lensing systems is carried out (Acevedo Barroso et al. 2025), finding 16 lens candidates (of which five have a convincing lens model) and a parallel search of strong lenses via the use of neural networks (Pearce-Casey et al. 2025) aimed at testing their performance on a real Euclid field. A similar search with neural networks in the other ERO fields is carried out in Nagam et al. (2025). The first statistically relevant sample of around $500$ strong lenses in Euclid has finally been collected with the data from the Euclid Quick Data Release 1 (Q1, Euclid Quick Release Q1 2025; Euclid Collaboration: Aussel et al. 2025), which also represents the first publicly available dataset of Euclid Wide Survey (EWS)-like images, processed with the same pipeline and reaching the same depth as the final survey. Despite covering only $63.1$ $\deg^{2}$, Q1 provides a powerful testbed for Euclid’s lens searching capabilities, with the strong lensing discovery engine at the forefront of these efforts (Euclid Collaboration: Walmsley et al. 2025; Euclid Collaboration: Rojas et al. 2025; Euclid Collaboration: Lines et al. 2025; Euclid Collaboration: Li et al. 2025; Euclid Collaboration: Holloway et al. 2025).
@@
High-resolution imaging, together with a large field of view, is crucial for detecting a large number of gravitational lenses. However, the substantial amount of data coming from the mission will necessarily require automated methods for both lens finding and lens modelling. Convolutional neural networks (CNNs) are already used extensively in lens-finding tasks, and generally perform well (Petrillo et al. 2017, 2019a, 2019b; Metcalf et al. 2019; Cañameras et al. 2020; Li et al. 2020; Rezaei et al. 2022; Rojas et al. 2022; Nagam et al. 2024). However, modelling of gravitational lenses has traditionally relied on complex and time-consuming techniques that require manual inputs, such as maximum likelihood and Markov chain Monte Carlo methods (Keeton 2016; Nightingale et al. 2021a, b; Harvey-Hawes & Wiltshire 2024). The advent of machine learning has allowed for the automation and acceleration of this process. In particular, CNNs and Bayesian neural networks (BNNs) have shown remarkable results in recovering lens parameters (Perreault Levasseur et al. 2017; Pearson et al. 2019, 2021; Schuldt et al. 2021, 2023b; Gentile et al. 2023). A recent work by Park et al. (2021) demonstrated that BNNs can successfully recover the posterior probability density functions of lens model parameters with sufficient accuracy for time delay cosmography, achieving sub-percent precision on the Hubble constant when applied to ensembles of lensed quasars. Complementary approaches using simulation-based inference, such as neural posterior estimation, have shown even better uncertainty calibration and accuracy compared to BNNs for strong lens modelling on ground-based data (Poh et al. 2022). Furthermore, Legin et al. (2022) introduced an efficient hierarchical inference framework that corrects for selection biases in strong lensing surveys using neural networks, allowing for fast and unbiased inference of population-level lens parameters.
@@
The main algorithm in this work is LEns MOdelling with Neural networks (LEMON; Gentile et al. 2023), a BNN able to perform fast automated analysis of strong gravitational lenses. In the original work, Gentile et al. (2023) used LEMON to estimate three parameters of a strong gravitational lens modelled as a singular isothermal ellipsoid (SIE), the Einstein radius and the two components of the ellipticity, for mock Hubble Space Telescope (HST) and Euclid gravitational lenses. Along with the expected values, LEMON is also able to output uncertainties for each parameter.
@@
In this work, we expand the results of Gentile et al. (2023), by predicting both mass and light parameters of the foreground galaxies in isolated galaxy-galaxy gravitational lenses, considering a single Srsic model for the light profile and using mock Euclid lenses with contaminants in the image for both training and testing. We also verify the behaviour of both uncertainty components and verify the capability of LEMON to generalise to real Euclidised HST lens images.
@@
The paper is structured as follows. In Sect. 2 we give details on the various datasets used for the analysis. In Sect. 3 we briefly introduce the LEMON algorithm. In Sect. 4 we describe the training procedure, the metrics used to quantify the performance of LEMON, and the calibration procedure for the total uncertainty. Section 5 is dedicated to the results of the analysis on the simulated test set, while Sect. 6 shows the results associated with real Euclidised lenses and real Euclid lenses found in the Perseus ERO and Q1 fields. Section 7 shows how the joint use of LEMON and classical modelling methods could provide both a speed up of the Euclid modelling pipeline and a reduction of its failure late. We finally give our conclusions in Sect. 8.
@@ 2 Data
In the following, we give details on the various datasets used for the analysis. In Sect. 2.1, the Euclid mock lenses used for training and testing are described. We then describe the real Euclidised lenses used for testing LEMON performance on real lenses in Sect. 2.2.
@@ 2.1 Euclid mock lenses
Table 1: Parameter ranges for the Euclid mock lenses. [^1] Only the Srsic index was extracted from a uniform distribution. This choice was made to avoid introducing an informative prior for this parameter. All the other parameters were extracted from non-analytic, empirical distributions derived from the Flagship simulation (Euclid Collaboration: Castander et al. 2025). The coverage of the parameters in the reported intervals is thus non-uniform. Notice that the intervals for the two ellipticity components are the same for both mass and light distributions, and as such have been reported together.
@@
$x$-ellipticity $\epsilon_{x}$ [0.80,0.80][-0.80,0.80]
$y$-ellipticity $\epsilon_{y}$ [0.80,0.80][-0.80,0.80]
@@
Lens Srsic index $n_{\mathrm{lens}}$ $[2.00,8.00]$
Lens magnitude $m_{\mathrm{lens}}$ $[13.95,23.56]$
Source $x$ position $x_{\textrm{src}}/\textrm{arcsec}$ [5.00,4.00][-5.00,4.00]
Source $y$ position $y_{\textrm{src}}/\textrm{arcsec}$ [4.00,8.00][-4.00,8.00]
@@
Figure 1: Distribution of mass and light parameters for the Euclid mock lenses.
For training and performance testing of LEMON, we used two simulation sets.
  • The first, or main dataset, consists of 100 000100\,000 gravitational lenses, modelled with a SIE mass profile plus external shear, with $\gamma_{x,y}\in[-0.03,0.03]$. In this set, the mass and light distributions are completely unaligned.
  • The second dataset is identical to the main dataset, but it allows for a wider range of external shear components, $\gamma_{x,y}\in[-0.4,0.4]$. This dataset was used to test LEMON’s ability to recover the shear components. The network trained with this dataset shows performances comparable to the main network we used in the main text, while also managing to recover the shear parameters. More information are available in Appendix A.
@@
All the datasets are simulated by the Strong Lensing Science Working Group of the Euclid Consortium (Metcalf et al., in prep.). The lens images were made by simulating the Euclid visible instrument (VIS, Euclid Collaboration: Cropper et al. 2025) camera output, at the depth of the EWS. The samples were selected to capture the full range of environmental scenarios that may occur when observing a gravitational lens, including objects along the line of sight as well as companion or background objects surrounding the lens system.
@@
The lenses were constructed by first considering galaxies in the Euclid Flagship simulation (Euclid Collaboration: Castander et al. 2025), taken from a sky area of the order of $100\,\textrm{deg}^{2}$, under the requirements that the galaxy chosen is brighter than the lens galaxy magnitude limit of 24 in the $I_{\scriptscriptstyle\rm E}$ band, which is the VIS imaging band, and categorised as a central galaxy. Central galaxies are galaxies that are positioned at the centres of the dark matter halos. Large halos also contain satellite galaxies, which should not produce isolated lenses like the ones we wish to model here. For each selected candidate, the apparent magnitude, effective radius, position angle, axis ratio and redshift were taken from the Flagship and converted into a surface brightness model. In general, the light model is composed of a disc and a spheroid component, but in our catalogue, only galaxies dominated by the spheroidal component are selected. This spheroid component is represented by a single Srsic profile. The total flux and effective radius is taken from the Flagship simulation. The Srsic index is uniformly distributed between 2 and 8. A mass model is then chosen by considering analytical profiles centred on each galaxy.
@@
The mass in the lens was normalised with the ratio of dark matter to stellar mass within one effective radius. The stellar mass and effective radius were taken from the Flagship simulation. The dark matter fraction within one effective radius was taken to be normally distributed with a mean of 0.6 and a standard deviation of 0.1, consistent with observations (Mukherjee et al. 2022).
@@
Once the lens galaxies had been selected, a source galaxy was placed in the relative background. This increases the sample size of lenses in the dataset, in order to offset the inherent rarity of ‘natural lenses’ (i.e. lenses in the Flagship simulation that naturally have a source behind them when the image is traced onto the observer plane). The redshift of the source was sampled randomly from the following distribution:
@@
where $\gamma=-0.23$ and $z_{0}=3.06$ are values found by fitting to the COSMOS 2020 Farmer catalogue $I$-band number counts (Weaver et al. 2022). The surface brightness is represented by between one and four Srsic profiles. One profile has a brightness and size based on randomly selected sources from the Hubble Ultra Deep Field (HUDF) with similar redshifts (see Meneghetti et al. 2008, 2010 for a discussion of this sample). Sixty percent of the time additional smaller components are added. They are positioned according to a Gaussian distribution centred on the main source with a scale size equal to the scale size of the host source. Each component is given a random orientation, and axis ratio. Their fluxes are drawn from a power-law distribution with index 4. This is a simplistic attempt to emulate some of the complex morphologies of high redshift sources.
Ray-shooting through the light cones was performed, to construct an image of the lens. The deflection caused by other objects along the line of sight is not included, so that the mass distribution remains a SIE in all cases. The lens was then rejected or accepted based on an observability criterion. This criterion seeks to filter out the many lenses candidates that would not be classified as lenses. The procedure is as follows. The lens light was subtracted from the lensed source light. Pixels that were greater than 4 times the background noise were considered to be in a source image. This region was divided into disconnected images. Images with a signal-to-noise (S/N) of less than 10 were discarded. If there was more than one image remaining, the object was classified as a lens. If there was one image, it could be an arc or a ring. If there was a hole in the single image it was considered a lens. Finally, if the image was crossed by a critical curve it was classified as a lens. From the visual inspection of many simulated lenses, it was found that this criterion accepts almost all the lenses that an inspector would identify as a lens.
A summary of the ranges of parameters for the mock lenses is listed in Table 1, with the respective distributions shown in Fig. 1. We underline the fact that, except for the Srsic index, all the parameters are drawn from non-analytic, empirical distributions derived from the Flagship simulation, so the coverage for these parameters is not uniform. An example of some mock Euclid lens images used for training is shown in Fig. 2.
The eight parameters listed in Table 1 are the main ones that we want to predict with LEMON. The recovery of a ninth parameter, the Einstein mass, is discussed in Appendix B. They are listed as follows.
@@
    where $A_{\textrm{Ein}}$ is the Einstein area in square radians, defined as the area of the largest critical curve, and the factor 648 000648\,000 corresponds to converting $\pi$ radians into arcseconds. Because $A_{\textrm{Ein}}$ is calculated numerically on a fixed resolution grid, it takes discrete values separated by 0.054 square arcseconds.
@@
    Ellipticity $x$ component, defined as
    $\epsilon_{x}=\frac{1-q}{1+q}\cos(2\phi)\,$, (3)
    where $q=b/a$ is the axis ratio of the mass or light distribution, while $b$ and $a$ are, respectively, the semi-minor and semi-major axes of the respective ellipsoidal distribution. The angle $\phi$ is, instead, the position angle of the mass or light distribution, which indicates the angle between the semi-major axis of the ellipsoidal distribution and the positive $x$ axis of the standard Cartesian reference frame.
@@
    Ellipticity $y$ component, defined as
    $\epsilon_{y}=\frac{1-q}{1+q}\sin(2\phi)\,$, (4)
    where $q$ and $\phi$ are defined as above.
@@
    Circularised effective radius of the lens, $R_{\textrm{e}}$, i.e. the radius enclosing one-half of the lens galaxy light, obtained from the major axis effective radius, $R_{\textrm{e, maj}}$, via the formula $R_{\textrm{e}}=\sqrt{q}\,R_{\textrm{e, maj}}$. We refer to it as $R_{\textrm{e, lens}}$ in the following.
@@
    Lens Srsic index, which regulates the slope of the surface brightness profile. It corresponds to the parameter $n$ that appears in the Srsic profile definition
@@
    where $I_{\textrm{e}}=I(R_{\textrm{e}})$ and $b_{n}$ is a parameter, solution to the equation $\Gamma(2n)=2\,\gamma(2n,b_{n})$, where $\Gamma$ and $\gamma$ are the complete and incomplete gamma functions, respectively. We refer to it as $n_{\textrm{lens}}$ in the following.
@@
    Lens magnitude, defined as the apparent magnitude in the $I_{\scriptscriptstyle\rm E}$ band of the main lens. It should be noted that, due to the fact that input images are based on counts instead of fluxes, we need to take into account zero-point shifts when predicting magnitudes of non-simulated lenses.
  7. 7.
    The $x$ and $y$ position of the source (in arcseconds), defined as the unlensed $x$ and $y$ co-ordinates of the sources brightest component, measured relative to the image centre.
The original simulated images have dimensions of $20\arcsec\times 20\arcsec$, as is shown in Fig. 2, but during the training of LEMON we create $10\arcsec\times 10\arcsec$ cut-outs of these images, centred on the lenses, in order to better see the lensing features.
Figure 2: $20\arcsec\times 20\arcsec$ original VIS images of some examples of Euclid mock lenses used for training LEMON.